OpenStudy (anonymous):

Solve for x: 3/2 (x-5) I know the answer is x=1, but how do you get that? I don't understand how to solve problems like these...

OpenStudy (anonymous):

where is the = sign

OpenStudy (amistre64):

your missing parts of the problem

OpenStudy (turingtest):

If it doesn't start as an equation it cannot become one. Perhaps you meant (3/2)(x-5)=-6 ?

OpenStudy (anonymous):

\[3/2(x-5)=-6\]

OpenStudy (anonymous):

Yes TurningTest, that is what meant. I'm sorry

OpenStudy (turingtest):

First get rid of the fraction. To do this multiply the whole equation by it's inverse. The inverse of 3/2=2/3, so we multiply BOTH SIDES by 2/3: (2/3)(3/2)(x-5)=-6(2/3) Make sure to multiply both sides!!! x-5=-4 now to get x by itself add 5 to BOTH SIDES x-5+5=-4+5 x=1